# Science Opportunities of Wet Extreme Mass-Ratio Inspirals

### v1

@@ I Introduction
Extreme mass-ratio inspirals (EMRIs) and massive black hole (BH) binaries are the two main extragalactic sources of space-borne gravitational wave detectors, such as LISA (Laser Interferometer Space Antenna) [1, 2, 3, 4], Tianqin [5, 6, 7, 8] and Taiji [8, 9, 10], which operate in the millihertz frequency band. In particular, EMRIs are known for their superior precision in measuring weak environmental forces, including tidal gravitational forces from nearby stellar-mass objects [11], disk forces associated with Active Galactic Nuclei (AGN) [12, 13, 14, 15, 16], and dynamical friction from clouds of Axion-like particles populated around massive BHs [17, 18, 19, 20].
Those EMRIs formed through the interaction with accretion disks are often referred as “Wet EMRIs” [21]. It was previously pointed out in [22] that star formation and evolution within AGN may lead to this class of objects. More recently, quantitative modeling of the nuclear star cluster and the accretion disk have shown that disk capture and migration of a stellar mass object may also give rise to rapid formation of wet EMRIs that eventually merge with the central massive BHs [21]. Based on a class of the disk and cluster models, the inferred rate is comparable to or larger than those EMRIs formed through gravitational scattering within the nuclear cluster (the “Dry EMRIs”) [23, 24].
@@ II.1 Population estimation
Table 1: Forecasted rates for stellar-mass black hole (sBH) EMRIs in the redshift range $4.50<z<$ 4.5, with the mass of sBH $\mu=10M_{\odot}$. The table presents total EMRI rates, LISA detection rates (with SNR20absent20$\geq$ 20≥ 20) and rates for which the host AGN galaxy is resolvable. Wet EMRI models follow the configuration of our previous work [21, 23], assuming a conservative AGN fraction $f_{\rm AGN}=1\%$ throughout the universe. For comparison, the last row shows dry EMRI rates with $N_{p}=10$ plunges per EMRI.
@@
The structure and dynamics of the accretion disk also influence the migration and formation rates of EMRIs. In $\alpha$-type disks, the viscosity scales with the total pressure (which includes both the gas and radiation pressure), while in $\beta$-type disks the viscosity scales with the surface density of the disk. These differences affect the disk’s structure and, consequently, the migration timescales and efficiency. Migration continues until the sBH approaches the SMBH, where GW emission dominates, driving the final inspiral. The presence of accretion flows accelerates both the capture of sBHs into the disk and their inward migration, significantly increasing the EMRI formation rate compared to the dry channel, where no disk is present [23].
@@
with a power index $\beta\approx 0$. This updated mass function offers a more accurate description of SMBHs in the mass range $10^{4}\sim 10^{7}$, which is critical for EMRI formation. Unlike earlier models that extrapolated SMBH populations from higher masses ($\sim 10^{7}M_{\odot}$[2], the TDE-based estimates are directly informed by observations of SMBHs with masses as low as $10^{5}M_{\odot}$ [44].
@@
Table 1 summarizes the updated wet EMRI population estimates based on the revised SMBH mass function. The table provides the total wet EMRI rates for sBHs ($\mu=10M_{\odot}$) over the redshift range $4.50<z<$ 4.5, as well as the corresponding LISA detection rates (for a signal-to-noise ratio $\mathrm{SNR}\geq 20$) and the fraction of resolvable host AGN galaxies. A host galaxy is considered resolvable if the expected number of AGNs within the localization volume is less than 1.1 (as detailed in Sec. II.5). These calculations are based on the framework established in our previous work [21, 23], assuming a conservative AGN fraction of $f_{\rm AGN}=1\%$ throughout the universe. For comparison, the final row includes the dry EMRI rates, assuming $N_{p}=10$, where $N_{p}$ represents the number of plunges per EMRI.
To access the LISA detection rates, we adopt the Augmented Analytic Kludge (AAK) waveform model [46, 47], which has been implemented in FEW [47, 48, 49] ). These waveforms are generated for the last two years of inspiral before coalescence and are expressed in the time domain for both GW polarizations. The LISA detector’s response to these waveforms is modeled using its two orthogonal channels (I and II), with the GW signals projected as [50, 51]
@@
[…]
@@ II.2 Transient Electromagnetic Counterparts: type II QPEs
In a gas-rich environment, it is natural to expect EM emissions as the sBH travels through the gas around the SMBH. If the sBH is aligned with and embedded in the gas disk, the continuous EM emission arising from gas accretion into the sBH will be reprocessed by the disk and will be hard to distinguish from the AGN background emission. In this case, the environmental effects on the gravitational waveform of EMRI have been investigated in previous studies [12, 13, 14, 15, 16]. Notice that the disk orientation at larger radii may not be the same as the local disk direction, where the local disk is likely perpendicular to the SMBH spin direction [54], while the outer disk direction depends on the angular momentum of the gas feeding at larger radii. The wrapped disk will exert a torque onto the central BH and align or antialign the BH spin with the angular momentum of the outer disk. The wrap radius is approximately [55, 56]
@@
and the disk alignment timescale in the small wrap case is approximately [57, 55, 58, 56, 59]
@@
where $\alpha$ is the standard viscosity coefficient in thin disks, $a$ is the dimensionless spin of the SMBH, $\dot{M}_{\bullet}$ is the SMBH accretion rate and the Eddington accretion rate $\dot{M}_{\bullet,\rm Edd}=M_{\bullet}/(5\times 10^{7}{\rm yr})$, and $h$ is the disk aspect ratio. If the accretion of the SMBH is coherent in a period much longer than $t_{\rm align}$, we expect an equatorial accretion disk and an equatorial EMRI when the EMRI enters the LISA band; therefore, no detectable EM transient is produced. Otherwise, if the gas feeding direction is random in each accretion episode with an accretion time scale shorter than $t_{\rm align}$, EM transients arising from EMRI and disk collisions are possible: for a sBH embedded in the disk, the Lense-Thirring precession tends to drive its orbit out of the disk plane, while the disk interaction tends to keep the sBH within the plane. We denote the critical radius as $r_{\rm dec}$ where the alignment timescale for the sBH orbit with the disk is equal to the Lense-Thirring precession timescale of the sBH. At smaller radii $r<r_{\rm dec}$, the sBH orbiting around the SMBH may hit the accretion disk and produce (quasi-)periodic EM transients, similar to some of the discussion of X-ray quasi-periodic eruptions (QPEs) [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], where eruptions are highly likely due to collisions between a low-frequency ($f_{\rm obt}\sim 10^{-5}$ Hz) EMRI and an accretion disk formed in a recent TDE [60, 61, 62, 63, 64, 34, 33, 65, 66, 67, 68, 69, 70]. For clarity, we will call the TDE-associated QPEs as type I QPEs and the quasi-periodic EM transients produced by an EMRI crossing an AGN disk as type II QPEs:
@@
where $\vec{a}$ is the dimensionless spin vector of the central SMBH and $r$ is the orbital radius. The disk interaction tends to align the secondary massive object (SMO) orbital plane with the disk plane at a rate $\vec{\omega}_{\rm DI}$ (in the direction of $\vec{n}_{\rm disk}\times\vec{L}$), where $\vec{n}_{\rm disk}$ is the normal vector of the disk and $\vec{L}$ is the angular momentum of the SMO orbit. The orbital angular momentum $\vec{L}$ evolves according to
@@
It is straightforward to see $\omega_{\rm DI}$ reaches it maximum value
@@
at $\iota_{\rm sd}\approx 2h$. The decoupling radius is determined by the equation $\omega_{\rm DI}(r_{\rm dec})=\omega_{\rm DI}^{\rm max}(r_{\rm dec})$ as
@@
for $\beta$ disks, where we have defined $\mu_{30}=\mu/30M_{\odot},M_{\bullet,6}=M_{\bullet}/10^{6}M_{\odot},\dot{M}_{\bullet,0.1}=\dot{M}_{\bullet}/0.1\dot{M}_{\bullet,\rm Edd}$ and we have used analytic models of $\alpha$ and $\beta$ disks [12, 13]. The two numerical factors are nearly equal because the two disk models are similar at large radii where the gas pressure dominates the total pressure in the accretion disk. Due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$, the decoupling radius $r_{\rm dec}$ turns out to be insensitive to $M_{\bullet},\alpha,m$, and only has a mild dependence on the SMBH accretion rate $\dot{M}_{\bullet}$. From Eq. (11), it is straightforward to see that the SMO is nearly aligned with the disk ($\vec{L}/L\approx\vec{n}_{\rm disk}$) for $r>r_{\rm dec}$ and starts to precess around the SMBH spin axis at $r<r_{\rm dec}$. The transition between the two regimes is very sharp due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$.
As an sBH crosses an accretion disk with relative velocity $v_{\rm rel}$ higher than the local gas sound speed $v_{\rm s}$, the gas within the accretion radius $r_{\rm acc}:=2Gm/v_{\rm rel}^{2}$ will be shocked and the total orbital energy loss of the sBH per collision turns out to be [65]
$\delta E_{\rm sBH}$ $=4\pi\ln\Lambda\frac{G^{2}m^{2}}{v_{\rm rel}^{2}}\frac{\Sigma}{\sin(\iota_{\rm sd})}\ ,$ (17)
$\approx 2\times 10^{46}{\rm ergs}\left(\frac{\ln\Lambda}{10}\right)\Sigma_{5}\mu_{30}^{2}r_{100}\left(\frac{\sin\iota_{\rm sd}}{0.1}\right)^{-3}\ ,$
@@
$\Sigma(r)=1.7\times 10^{5}\ {\rm g\ cm}^{-2}\alpha_{0.01}^{-1}\dot{M}_{\bullet,0.1}^{-1}r_{100}^{3/2}\ ,$ (18)
@@
As the EMRI enters the LISA sensitivity band with orbital radius $r\sim 10M_{\bullet}$ and orbital frequency $f_{\rm obt}\sim$ mHz, the energy budgets per collision for two disk models are quite different. In the $\alpha$-disk model, it scales as $\delta E_{\rm sBH}\propto r^{5/2}\propto f_{\rm obt}^{-5/3}$, that is say, the mHz QPEs will be $\mathcal{O}(10^{3})$ times weaker than their low-frequency counterparts. Although the identification of mHz type II QPEs is difficult in a sea of much stronger AGN variability, the joint GW measurement of the orbits may help identify the detection of mHz type II QPEs. If the accretion disk can be modeled as a $\beta$-disk, the energy budget per collision scales as $\delta E_{\rm sBH}\propto r^{2/5}\propto f_{\rm obt}^{-4/15}$, i.e. the mHz QPEs will be a few times weaker than their low-frequency counterparts. In this case, the identification of mHz QPEs is possible even without the input of GW observations. This multi-messenger source will be a convenient probe to accretion physics, e.g., distinguishing the two disk models and inferring the magnitude of the commonly used but not well constrained $\alpha$ parameter.
@@
To summarize, wet EMRIs are promising sources of EM transients (dubbed as type II QPEs). If confirmed, they can be used as a probe to AGN disk structure, and their existence provides indirect evidence for warped AGN disks and chaotic AGN accretion. In addition, wet EMRIs in the mHz band will be emitting multi-messenger signals that are possibly detectable by both spaceborne GW detectors and X-ray telescopes, where there is a chance that the X-ray flares get lensed by the SMBH.

### v2

@@ I Introduction
Extreme mass-ratio inspirals (EMRIs) and massive black hole (BH) binaries are the two main extragalactic sources of space-borne gravitational wave detectors, such as LISA (Laser Interferometer Space Antenna) [1, 2, 3, 4, 5], TianQin [6, 7, 8, 9] and Taiji [9, 10, 11], which operate in the millihertz frequency band. In particular, EMRIs are known for their superior precision in measuring weak environmental forces, including tidal gravitational forces from nearby stellar-mass objects [12], disk forces associated with Active Galactic Nuclei (AGN) [13, 14, 15, 16, 17], and dynamical friction from clouds of Axion-like particles populated around massive BHs [18, 19, 20, 21].
Those EMRIs formed through the interaction with accretion disks are often referred as “Wet EMRIs” [22]. It was previously pointed out in [23] that star formation and evolution within AGN may lead to this class of objects. In recent years, there has been growing interest in the environmental effects influencing EMRI formation [24, 25, 26, 27, 28, 29, 30]. More recently, quantitative modeling of the nuclear star cluster and the accretion disk have shown that disk capture and migration of a stellar mass object may also give rise to rapid formation of wet EMRIs that eventually merge with the central massive BHs [22]. Based on a class of the disk and cluster models, the inferred rate is comparable to or larger than those EMRIs formed through gravitational scattering within the nuclear cluster (the “Dry EMRIs”) [31, 32].
@@ II.1 Population estimation
[…]
@@
The structure and dynamics of the accretion disk also influence the migration and formation rates of EMRIs. In $\alpha$-type disks, the viscosity scales with the total pressure (which includes both the gas and radiation pressure), while in $\beta$-type disks the viscosity scales with the surface density of the disk. These differences affect the disk’s structure and, consequently, the migration timescales and efficiency. Migration continues until the sBH approaches the SMBH, where GW emission dominates, driving the final inspiral, as illustrated in FIG. 1. The presence of accretion flows accelerates both the capture of sBHs into the disk and their inward migration, significantly increasing the EMRI formation rate compared to the dry channel, where no disk is present [31].
@@
with a power index $\beta$ set to zero in the simulation. This updated mass function provides an empirical measurement of SMBH distribution, based on recent TDE observations [52]. Unlike other models—such as the Population III seed model, which is derived from simulations, and the phenomenological model, which is constructed from theoretical assumptions—the TDE-based mass function is directly informed by observational data, making it potentially more reliable.
In contrast to earlier studies that extrapolated SMBH populations from higher masses ($\sim 10^{7}M_{\odot}$[2], the TDE-based estimates are grounded in observations of SMBHs with masses as low as $10^{5}M_{\odot}$ [52], making them particularly relevant for modeling EMRI sources in the LISA band. Although current TDE observations are limited to the local universe ($z<0.5$), we assume this mass function to be universal and extrapolate its shape to higher redshifts ($z\sim 4.5$). This extrapolation is motivated by the assumption that the fundamental black hole demographics and EMRI formation mechanisms do not evolve significantly over this redshift range.
II.1.1 Population and detectability
@@
Table 1 summarizes the updated wet EMRI population estimates based on the revised SMBH mass function. The table provides the total wet EMRI rates for sBHs ($\mu=10M_{\odot}$) over the redshift range $0<z<4.5$, as well as the corresponding LISA detection rates (for a signal-to-noise ratio $\mathrm{SNR}\geq 20$) and the fraction of resolvable host AGN galaxies. A host galaxy is considered resolvable if the expected number of AGNs within the localization volume is less than 1.1 (as detailed in Sec. II.5). These calculations are based on the framework established in our previous work [22, 31], assuming a conservative AGN fraction of $f_{\rm AGN}=1\%$ throughout the universe. For comparison, the final row includes the dry EMRI rates, assuming $N_{p}=10$, where $N_{p}$ represents the number of plunges per EMRI.
To evaluate LISA detection rates and conduct the Fisher matrix analyses presented in this work, we adopt the Augmented Analytic Kludge (AAK) waveform model [54, 55, 56, 57, 58, 59], as implemented in Fast EMRI Waveforms (FEW) [59, 60, 61, 62, 63]. All analyses are carried out in a 14-dimensional parameter space that includes the mass and spin of the SMBH, the mass of the sBH, the luminosity distance, the initial semi-latus rectum, initial eccentricity, sky location, orbital inclination, and the initial orbital phases in $\phi,\theta,r$. For both dry and wet EMRIs, we assume isotropic distributions for sky location and SMBH spin direction. The three phase angles are sampled uniformly from $0,2\pi]$, coalescence times are randomly drawn from the interval $0,2\pi]$ years, and the cosine of the inclination angle is sampled uniformly from [1,1][-1,1], The initial eccentricity $e_{0}$ is assigned based on the formation channel: for dry EMRIs, $e_{0}$ is drawn from a uniform distribution in $[0,0.2]$, reflecting their dynamical origins. For wet EMRIs, we assume circular orbits with $e_{0}=0$, consistent with strong eccentricity damping in AGN disks.
The LISA detector’s response to these waveforms is modeled using its two orthogonal channels (I and II), with the GW signals projected as [64, 65]
@@
$\langle a,b\rangle=4\,\text{Re}\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{\tilde{a}(f)\tilde{b}^{*}(f)}{P_{n}(f)}\,df.$ (7)
@@
II.1.2 Identifications of wet EMRIs and host AGNs
The distinction between wet and dry EMRIs can be drawn based on their orbital characteristics and potential environmental signatures in the GW signal.
In wet environments, based on standard migration theory, strong interactions between sBHs and the accretion disks of rapidly accreting SMBHs—primarily via dynamical friction and disk-induced density waves—efficiently damp the vertical momentum and eccentricity of a sBH, aligning it with the disk plane and circularizing its orbit. Consequently, wet EMRIs typically exhibit near-zero eccentricities and low inclinations by the time they enter the LISA band.
In contrast, dry EMRIs, which form through dynamical scattering in gas-poor nuclear star clusters, where such dissipative effects are absent. These systems typically retain high eccentricities and inclined orbits throughout their evolution [22, 31]. The capability of space-based GW detectors such as LISA to measure orbital eccentricities with high precision—for instance, down to $e\sim 10^{-5}$ [53, 68]—enables the potential classification of EMRIs based on their orbital morphology.
Beyond orbital shape, wet EMRIs may also be distinguished through other observational features. One such possibility arises from the influence of the AGN disk on the sBH’s orbital evolution. If the disk is sufficiently dense, the orbiting sBH can experience torques or perturbations from the surrounding gas. These interactions may imprint detectable features on the GW waveform, such as phase shifts or resonant modulations, offering direct evidence of a disk-assisted inspiral [69, 70]. Another avenue for distinction involves EM counterparts. Wet EMRIs may generate Type II QPEs — transient X-ray flares that occur when the sBH crosses the warped AGN disk twice per orbit (see Sec. II.2). The contemporaneous detection of such QPEs alongside the GW signal would provide compelling multi-messenger confirmation of a wet EMRI event. Even in the absence of a direct EM counterpart, the spatial association of an EMRI signal with a known AGN within the GW localization volume increases the probability that the event originated in a gas-rich environment.
To evaluate the potential for identifying the AGN host of an EMRI event, we utilize Fisher matrix analysis, a well-established statistical method to estimate the precision of parameter measurements from observed signals. The Fisher information matrix, $\Gamma_{ij}$, is defined as
where $h$ is the signal, $\lambda_{i}$ are the source parameters, and the inner product is defined by Eqn. 7. The covariance matrix, $\Sigma_{ij}=(\Gamma^{-1})_{ij}$, is obtained by inverting the Fisher matrix, with the diagonal elements, $\sigma_{\lambda_{i}}^{2}=\Sigma_{ii}$, represent the variances of the individual parameters, thereby quantifying the uncertainties and correlations associated with the source localization, including angular resolution.
To compute the localization volume—the three-dimensional region in which the EMRI source is likely located—we use catalogs of wet EMRIs generated from a range of astrophysical models (see Table 1). These include, for example, an $\alpha$-disk model with a characteristic disk lifetime $T_{\rm disk}=10^{6}$ years. For each model, we estimate the number of potential AGN hosts located within the expected localization volumes.
The solid angle uncertainty, $\Delta\Omega_{s}$, is derived from the angular components of the Fisher matrix, associated with the angular parameters $(\theta_{s},\phi_{s})$, and is given by (see also [71, 72])
Combining this with the uncertainty in luminosity distance $d_{L}$, the three-dimensional sky localization volume is calculated as
$V_{\text{sky}}=\frac{1}{2}r^{3}(z)\,\sigma(\ln{d_{L}})\,\Delta\Omega_{s}\,$, (10)
where $r(z)$ is the comoving distance to the source at redshift $z$, and $\sigma(\ln{d_{L}})$ is the relative uncertainty in the luminosity distance measurement. For EMRIs detected by LISA, Fisher analysis indicates that $V_{\text{sky}}$ can be restricted to $\mathcal{O}(10^{2})\,\rm Mpc^{3}$ at moderate redshifts ($z<0.3$ [71]), depending on the SNR and the inclusion of higher order waveform harmonics, demonstrating the high localization precision of LISA.
To assess whether an EMRI can be classified as a bright siren, we adopt an average number of galaxies per comoving volume $\bar{n}_{\rm gal}=10^{-2}\rm Mpc^{-3}$ [73] and assume a conservative AGN fraction of $f_{\rm AGN}=1\%$. Although observational studies suggest that the AGN fraction is approximately $1\%$ at low redshift and may increases to $1-10\%$ at higher redshifts, we adopt a constant, redshift-independent AGN fraction of $1\%$ throughout this work to ensure a conservative and robust estimate.
A system is defined as a bright siren if the expected number of AGNs within the localization volume is fewer than $1.1$. Across the various astrophysical models listed in Table 1, we find that approximately $20\%\sim 25\%$ of detectable wet EMRIs meet this criterion. This corresponds to an annual detection rate of 3–30 bright sirens, depending on the model and detection assumptions.
In contrast, systems with multiple AGN candidates within their localization volumes are classified as dark sirens, reflecting the ambiguity in host identification. To account for observational limitations—such as interstellar extinction and incomplete AGN catalogs — we conservatively assume that $10\%$ of initially identified bright sirens should be reclassified as dark sirens.
@@ II.2 Transient Electromagnetic Counterparts: type II QPEs
In a gas-rich environment, it is natural to expect EM emissions as the sBH travels through the gas around the SMBH. If the sBH is aligned with and embedded in the gas disk, the continuous EM emission arising from gas accretion into the sBH will be reprocessed by the disk and will be hard to distinguish from the AGN background emission. In this case, the environmental effects on the gravitational waveform of EMRI have been investigated in previous studies [13, 14, 15, 16, 17]. Notice that the disk orientation at larger radii may not be the same as the local disk direction, where the local disk is likely perpendicular to the SMBH spin direction [74], while the outer disk direction depends on the angular momentum of the gas feeding at larger radii. The warpped disk will exert a torque onto the central BH and align or antialign the BH spin with the angular momentum of the outer disk. The warp radius is approximately [75, 76]
@@
and the disk alignment timescale in the small warp case is approximately [77, 75, 78, 76, 79]
@@
where $\alpha$ is the standard viscosity coefficient in thin disks, $a$ is the dimensionless spin of the SMBH, $\dot{M}_{\bullet}$ is the SMBH accretion rate and the Eddington accretion rate $\dot{M}_{\bullet,\rm Edd}=M_{\bullet}/(5\times 10^{7}{\rm yr})$, and $h$ is the disk aspect ratio. If the accretion of the SMBH is coherent in a period much longer than $t_{\rm align}$, we expect an equatorial accretion disk and an equatorial EMRI when the EMRI enters the LISA band; therefore, no detectable EM transient is produced. Otherwise, if the gas feeding direction is random in each accretion episode with an accretion time scale shorter than $t_{\rm align}$, EM transients arising from EMRI and disk collisions are possible: for a sBH embedded in the disk, the Lense-Thirring precession tends to drive its orbit out of the disk plane, while the disk interaction tends to keep the sBH within the plane. We denote the critical radius as $r_{\rm dec}$ (as illustrated in FIG. 1), where the alignment timescale for the sBH orbit with the disk is equal to the Lense-Thirring precession timescale of the sBH. At smaller radii $r<r_{\rm dec}$, the sBH orbiting around the SMBH may hit the accretion disk and produce (quasi-)periodic EM transients, similar to some of the discussion of X-ray quasi-periodic eruptions (QPEs) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], where eruptions are highly likely due to collisions between a low-frequency ($f_{\rm obt}\sim 10^{-5}$ Hz) EMRI and an accretion disk formed in a recent TDE [80, 81, 82, 83, 84, 42, 41, 85, 86, 87, 88, 89, 90]. For clarity, we will call the TDE-associated QPEs as type I QPEs and the quasi-periodic EM transients produced by an EMRI crossing an AGN disk as type II QPEs:
@@
Figure 1: Different effects act at different radial scales. $R_{\rm warp}$ marks the warp radius of the accretion disk, within which the disk is warped by the relativistic Lense-Thirring effect (frame-dragging), and remains flat beyond. For $r<r_{\rm dec}$, Lense-Thirring precession dominates the orbital evolution of the sBH, potentially driving it out of the disk plane and causing misalignment. When $r<R_{\rm GW}$, GW emission governs the sBH’s migration (GW-dominated regime), while for $r>R_{\rm GW}$, migration is dominated by disk effects, such as gas torques (disk-dominated regime).
@@
where $\vec{a}$ is the dimensionless spin vector of the central SMBH and $r$ is the orbital radius. The disk interaction tends to align the sBH’s orbital plane with the disk plane at a rate $\vec{\omega}_{\rm DI}$ (in the direction of $\vec{L}\times\vec{n}_{\rm disk}$), where $\vec{n}_{\rm disk}$ is the normal vector of the disk and $\vec{L}$ is the angular momentum of the sBH’s orbit. The orbital angular momentum $\vec{L}$ evolves according to
@@
It is straightforward to see $\omega_{\rm DI}$ reaches its maximum value
@@
at $\iota_{\rm sd}\approx 2h$. The decoupling radius is determined by the equation $\omega_{\rm LT}(r_{\rm dec})=\omega_{\rm DI}^{\rm max}(r_{\rm dec})$ as
@@
for $\beta$ disks, where we have defined $\mu_{30}=\mu/30M_{\odot},M_{\bullet,6}=M_{\bullet}/10^{6}M_{\odot},\dot{M}_{\bullet,0.1}=\dot{M}_{\bullet}/0.1\dot{M}_{\bullet,\rm Edd}$ and we have used analytic models of $\alpha$ and $\beta$ disks [13, 14]. The two numerical factors are nearly equal because the two disk models are similar at large radii where the gas pressure dominates the total pressure in the accretion disk. Due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$, the decoupling radius $r_{\rm dec}$ turns out to be insensitive to $M_{\bullet},\alpha,\mu$, and only has a mild dependence on the SMBH accretion rate $\dot{M}_{\bullet}$. From Eqn. (14), it is straightforward to see that the sBH is nearly aligned with the disk ($\vec{L}/L\approx\vec{n}_{\rm disk}$) for $r>r_{\rm dec}$ and starts to precess around the SMBH spin axis at $r<r_{\rm dec}$ (as illustrated in FIG. 1). The transition between the two regimes is very sharp due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$.
As an sBH crosses an accretion disk with relative velocity $v_{\rm rel}$ higher than the local gas sound speed $v_{\rm s}$, the gas within the accretion radius $r_{\rm acc}:=2G\mu/v_{\rm rel}^{2}$ will be shocked and the total orbital energy loss of the sBH per collision turns out to be [85]
$\delta E_{\rm sBH}$ $=4\pi\ln\Lambda\frac{G^{2}\mu^{2}}{v_{\rm rel}^{2}}\frac{\Sigma}{\sin(\iota_{\rm sd})}\ ,$ (20)
[…]
@@
$\Sigma(r)=1.7\times 10^{6}\ {\rm g\ cm}^{-2}\alpha_{0.1}^{-1}\dot{M}_{\bullet,0.1}^{-1}M_{\bullet,6}^{2}r_{100}^{3/2}\ ,$ (21)
@@
As the EMRI enters the LISA sensitivity band with orbital radius $r\sim 10M_{\bullet}$ and orbital frequency $f_{\rm obt}\sim$ mHz, the energy budgets per collision for two disk models are quite different. In the $\alpha$-disk model, it scales as $\delta E_{\rm sBH}\propto r^{5/2}\propto f_{\rm obt}^{-5/3}$, that is say, the mHz QPEs will be $\mathcal{O}(10^{3})$ times weaker than their low-frequency counterparts. Although the identification of mHz type II QPEs is difficult in a sea of much stronger AGN variability, the joint GW measurement of the orbits may help identify the detection of mHz type II QPEs. If the accretion disk can be modeled as a $\beta$-disk, the energy budget per collision scales as $\delta E_{\rm sBH}\propto r^{2/5}\propto f_{\rm obt}^{-4/15}$, i.e. the mHz QPEs will be a few times weaker than their low-frequency counterparts. In this case, the identification of mHz QPEs is possible even without the input of GW observations.
This multi-messenger source will be a convenient probe to the two disk models. The EMRI parameters, including the SMBH parameters $M_{\bullet},a_{\bullet}$, the stellar-mass BH mass $\mu$ and the EMRI orbital inclination angle $\iota_{\rm sd}$, are expected to be tightly constrained from the GW signal [2]. In combination with the AGN luminosity $L_{\bullet}$, it is straightforward to calculate the the SMBH accretion rate $\dot{M}_{\bullet}$. With these information, viscosity parameter $\alpha$ is the only unknown parameter in $\alpha$- or $\beta$-disk models, and the two disk models predict largely different disk surface densities at $r\lesssim 10M_{\bullet}$ with a ratio
$\frac{\Sigma_{\alpha}(r)}{\Sigma_{\beta}(r)}=2.7\times 10^{-2}\alpha_{0.1}^{-1/5}\dot{M}_{\bullet,0.1}^{-8/5}M_{\bullet,6}^{-1/5}r_{10}^{21/10}\ ,$ (23)
which is insensitive to the unknown parameter $\alpha$. Since the energy budget of EMRI-disk collisions is proportional to the disk surface density $\delta E_{\rm sBH}\propto\Sigma$, collisions with a $\beta$-disk naturally produce luminous QPEs that cannot be reconciled with the $\alpha$-disk prediction.
@@
To summarize, wet EMRIs are promising sources of EM transients (dubbed as type II QPEs). If confirmed, they can be used for distinguishing different accretion disk models, and their existence provides indirect evidence for warped AGN disks and chaotic AGN accretion. In addition, wet EMRIs in the mHz band will be emitting multi-messenger signals that are possibly detectable by both spaceborne GW detectors and X-ray telescopes, where there is a chance that the X-ray flares get lensed by the SMBH.